# Deep Time Series Models: A Comprehensive Survey and Benchmark

### v1

@@ Abstract
Time series, characterized by a sequence of data points arranged in a discrete-time order, are ubiquitous in real-world applications. Different from other modalities, time series present unique challenges due to their complex and dynamic nature, including the entanglement of nonlinear patterns and time-variant trends. Analyzing time series data is of great significance in real-world scenarios and has been widely studied over centuries. Recent years have witnessed remarkable breakthroughs in the time series community, with techniques shifting from traditional statistical methods to advanced deep learning models. In this paper, we delve into the design of deep time series models across various analysis tasks and review the existing literature from two perspectives: basic modules and model architectures. Further, we develop and release Time Series Library (TSLib) as a fair benchmark of deep time series models for diverse analysis tasks, which implements 24 mainstream models, covers 30 datasets from different domains, and supports five prevalent analysis tasks. Based on TSLib, we thoroughly evaluate 12 advanced deep time series models on different tasks. Empirical results indicate that models with specific structures are well-suited for distinct analytical tasks, which offers insights for research and adoption of deep time series models. Code is available at https://github.com/thuml/Time-Series-Library.
@@
Time series refers to a sequence of data points indexed in a discrete-time order [1, 2], which are omnipresent in real-world applications, such as financial risk assessment, energy sustainability, and weather forecasting. Driven by the increasing availability of vast amounts of time series data across various domains, the community of time series analysis has witnessed tremendous advancements. Compared to image and text data, which have objectively prescribed syntax or intuitive patterns, the semantic information of time series data is primarily derived from the temporal variation [3]. This presents significant challenges in understanding the data, such as identifying sequential dependencies, trends, seasonal patterns, and complicated dynamics. Consequently, analyzing time series data requires sophisticated methods to capture and utilize these complex temporal representations.
@@
Given the crucial role of time series data in real-world applications [4, 5, 6], time series analysis has been a longstanding research direction. Time series analysis encompasses the process of analyzing the temporal variation to understand time series data and make accurate predictions and informed decisions. One of the essential cornerstone of time series analysis is discovering the underlying patterns in time series data, which involves the intricate temporal dependencies and variate correlations inherent within the data. By capturing these complex dependencies, time series models can effectively reveal the underlying dynamics, and facilitate various downstream tasks, including forecasting, classification, imputation, and anomaly detection.
@@
Traditional time series methods, such as AutoRegressive Integrated Moving Average (ARIMA) [1], Exponential Smoothing, and Spectral Analysis [7], have long served as stalwart tools in time series analysis. These models, grounded in statistical methodologies, have been instrumental in discovering patterns, trends, and seasonality within temporal variations. However, their capabilities are hindered due to the inherent limitations of capturing complex nonlinear relationships and long-term dependencies present in real-world time series data. The rigid assumptions of linearity and stationarity that underpin traditional models constrain their adaptability to eventful and evolving data flows.
| Survey | Analysis Task | Model Architecture | Benchmark | |||||||
| Forecasting | Classification | Imputation | Anomaly | MLP | CNN | RNN | GNN | Transformer | ||
| Fawaz et al. (2019) [8] | ✓ | ✓ | ✓ | ✓ | ||||||
| Braei et al. (2020) [9] | ✓ | ✓ | ✓ | ✓ | ||||||
| Torres et al. (2021) [10] | ✓ | ✓ | ✓ | ✓ | ||||||
| García et al. (2021) [11] | ✓ | ✓ | ✓ | ✓ | ||||||
| Wen et al. (2022) [12] | ✓ | ✓ | ✓ | ✓ | ||||||
| Jin et al. (2023) [13] | ✓ | ✓ | ✓ | ✓ | ✓ | |||||
| Shao et al. (2023) [14] | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | |||
| Qiu et al. (2024) [15] | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | |||
| Our Survey | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ |
@@
Deep models have garnered significant attention and achieved remarkable performance across various domains, including natural language processing (NLP) [16], [17], computer vision (CV) [18], [19], and recommendation systems [20]. In recent years, deep learning models [21, 22, 23, 24, 3] have demonstrated their capability to capture the intricate dependencies within time series data, making deep learning models a powerful tool for time series analysis over traditional statistical methods. More recently, Transformer models with attention mechanisms, originally developed for natural language processing tasks, have presented stunning power in processing large-scale data [17] and have also been adapted for learning time series data. These architectures offer the advantage of selectively focusing on different parts of the input sequence, allowing for more nuanced discovery of both temporal and variable dependencies in time series.
@@
Related Surveys Although various time series models designed for different analysis tasks have emerged in recent years, there is a lack of a comprehensive overview of existing methods, covering both tasks and models. Previous reviews focus exclusively on either a specific model architecture or an analysis task. For example, [8, 9, 10, 11] reviews deep learning methods for specific time series analysis tasks while failing to include advanced architecture such as Transformer. Several surveys [12, 13] provide up-to-date reviews for time series analysis focusing on specific deep learning architectures(i.e., Graph Neural Network and Transformer). Recently, BasicTS [14] and TFB [15] introduce forecasting benchmarks that enable an unbiased evaluation of existing approaches but do not provide an overview of the architectural design of those deep models.
@@
In this survey, we provide a comprehensive review of deep time series models for researchers and practitioners, starting from the basic modules to modern architectures. To foster practical applications, a time series benchmark is offered for a fair evaluation and identifying the effective scope of existing models. Our survey is organized as follows. Section 2 provides the background concepts of time series analysis. Section 3 introduces the basic modules that are widely utilized in prevalent deep time series models. Section 4 reviews the existing deep time series models in terms of the architecture design. Section 5 introduces the proposed open-source benchmark—Time Series Library (TSLib)—and presents extensive experimental comparison with detailed analysis. Section 6 provides a brief discussion of future research directions while Section 7 summarizes this survey.
@@ 2.1 Time Series
Time series is a sequence of observations ordered by time, which can be denoted as , where represents the observed values at time point and is the number of variables. Since time series data are physical measurements obtained from sensors, systems are often recorded with multiple variables. Consequently, real-world time series usually recorded in a multivariate form. Theoretical studies [25, 26] have shown that when there are two or more non-stationary series, a linear combination of them can be stationary. This co-integration property helps in uncovering and modeling long-term relationships among non-stationary series. Therefore, the essence of time series analysis is to capture and utilize the temporal dependencies and inter-variable correlations within the observations.
Temporal Dependency Given the sequential nature inherent in the observations, one evident technological paradigm is to capture the temporal dependence of a set of historical data. The basic idea of temporal dependencies is the intricate correlations between time points or sub-series. Traditional statistical models have laid the groundwork for modeling temporal dependencies. Prominent models include ARIMA (Autoregressive Integrated Moving Average) [1] have been extensively studied for capturing complex temporal patterns in the time series modality. Owing to their simplicity and interpretability, these statistical methods remain popular for tasks where the underlying temporal dynamics do not exhibit high complexity. Considering the high-dimensionality and non-stationarity of real-world time series, the research focus shifted towards deep learning for time series analysis. These advanced methods are designed to handle more complex temporal dynamics and offer greater flexibility in capturing the temporal dependency of time series data.
Variate Correlation In addition to capturing temporal dependencies, understanding the variate correlations within high-dimensionality plays a pivotal role in analyzings multivariate time series. These correlations refer to the complex interactions and associations among different variables changing across the time. They provide valuable insights into the underlying dynamics and dependencies among the measurements, enabling a more comprehensive understanding of the latent process. Traditional approaches, such as Vector Autoregressive (VAR) models [27], extend the concept of autoregression to multiple variables and can capture the relationships between multiple quantities as they evolve over time. Technically, VAR represents each variable as a linear combination of its lagged values and the lagged values of all other variables in the model, which results in an inability to capture complex and non-linear relationships. Recently, advanced deep models, such as Graph Neural Networks [28] and Transformers [29, 30], have also been introduced for variate correlation modeling.
@@ 2.2 Time Series Analysis Tasks
Based on the understanding of underlying patterns and trends within time series data, time series analysis encompasses various downstream applications, including forecasting [31, 32], imputation [33, 34, 35], classification [8, 36], and anomaly detection [37, 5], each serving distinct purposes in diverse application domains.
We illustrate representative time series analysis tasks in Figure 1. Forecasting is a fundamental task in time series analysis that requires models to uncover temporal dependencies and dynamic patterns in the data. By capturing the relationships between past and future data, the forecasting model aims to predict future values or trends of the input series. Missing data due to sensor failures, data corruption, or absent measurements are ubiquitous in practical applications, leading to a growing demand for time series imputation to obtain higher-quality data. Unlike forecasting, which predicts future values based on historical observations, imputation focuses on reconstructing missing values using the available contextual information. Anomaly detection involves identifying unusual or abnormal patterns within a time series, which can indicate critical events, system faults, or outliers requiring further investigation. Lastly, classification assigns a label or category to a given time series based on its characteristics, a task widely utilized in fields such as medical diagnosis.
@@ 3 Basic Modules
Time series modeling approaches have evolved significantly, transitioning from traditional statistical models to sophisticated deep learning models. Despite these advancements, many classical tools and analytical algorithms remain widely used and continue to serve as foundational design principles in modern deep models. In this section, we focus on the major tools of classical time series analysis and demonstrate how they have been integrated as fundamental components in contemporary deep time series models.
@@ 3.1 Stationarization
As a foundational concept in time series analysis, stationarity refers to the property of a time series where its statistical properties remain constant over time. A stationary time series has a constant mean and variance, which simplifies statistical analysis and makes it easier to capture the underlying patterns and behavior within a time series. Since many statistics-based time series analysis methods take stationarity as a basic assumption, stationarization of time series data has become an essential module. There are ways of transforming non-stationary time series into stationary. Traditional time series models stationarize the time series through differencing or log-transformation. In recent deep learning approaches, data normalization [38] takes the role of stationarization in a simple but effective way, which standardizes the value distribution of observations while maintaining the intrinsic variations and further helps mitigate the distribution shift between the source and target domains.
The deep adaptive input normalization (DAIN) layer [39] was proposed to adaptively stationarize time series data according to their original distribution. RevIN [40] introduces reversible instance normalization to time series data, which is an effective normalization-and-denormalization method with learnable affine transforms to make the model bypass the non-stationary inputs. Non-Stationary Transformer [41] (Stationary for short in the following) proposes a simpler but more effective series stationarization technique that improves the predictive capability of non-stationary series without extra parameters. Specifically, for a sequence with time stamps and variates , the outline of Stationary [41] can be summarized as:
where is in a small value for numerical stability. are the variate-specific mean and variance. To recover the distribution and non-stationarity of the original series, a de-normalization module is further used to augment the model output with mean and variance statistics of inputs.
The idea of stationarization and the above-mentioned techniques have been widely used in subsequent deep time series models [23, 30, 42, 43]. Recent SAN [44] rethinks the nature of non-stationary data and tries to split it into non-overlap equally-sized slices and perform normalization on each slice. Specifically, based on the evolving trends of statistical properties, SAN introduces a statistics prediction module to predict the distributions of future slices.
@@ 3.2 Decomposition
Decomposition [45, 46], as a conventional approach in time series analysis, can disentangle time series into several components with categorized patterns, and works primarily useful for exploring complex series variations. In the previous work, diverse decomposition paradigms are explored.
@@ 3.2.1 Seasonal-Trend Decomposition
Seasonal-trend decomposition [47, 48] is one of the most common practices to make raw data more predictable, which can separate the series into several different components: trend, seasonal, cyclical, and irregular, namely
where the trend component represents the overall long-term pattern of the data over time, the cyclical component reflects repeated but non-periodic fluctuations within data, the seasonal component indicates the repetitive patterns over a fixed period, and the irregular component is the residuals or remainder of the time series after the other components have been removed.
The trend-seasonality decomposition can be achieved by using mathematical tools such as filters or exponential smoothing [49, 50]. Previous statistical approaches mainly adopt the trend-seasonality decomposition as data pre-processing [51]. In deep models, Autoformer [22] firstly introduces the idea of decomposition to deep learning architecture and proposes a series decomposition block as a basic module to extract the seasonal and trend-cyclical parts of deep features and input series, whose computation process can be formalized as:
The series decomposition block is concisely implemented based on a temporally average pooling layer with padding operation to keep the sequence length unchanged. This design can capture trends , and the remainder is taken as the seasonal part . The proposed series decomposition block has been widely used in the follow-up [6, 52, 53, 54, 55] as a native building block of deep models to disentangle the underlying patterns of deep features.
@@ 3.2.2 Basis Expansion
Basis expansion is a mathematical method used to represent a function or a set of data points in terms of a new set of pre-defined functions. These new functions form a basis for a function space, meaning any function in that space can be expressed as a linear combination of these basis functions. In the context of time series analysis, basis expansion is used to reveal complex non-linear temporal relationships by decomposing the time series into a combination of basic variations, which also enhances interpretability. As a representative model, N-BEATS [56] presents hierarchical decomposition to time series by utilizing a fully connected layer to produce expansion coefficients for both backward and forward forecasts. For -th blocks in the proposed hierarchical architecture, the operation can be as follows:
@@
Subsequently, N-HiTs [57] redefine the N-BEATS by incorporating subsampling layers before the fully connected blocks, which enhances the input decomposition via multi-frequency data sampling and future predictor via multi-scale interpolation. DEPTS [58] puts forward a novel decoupled formulation for periodic time series by introducing the periodic state as a hidden variable and then develops a deep expansion module on top of residual learning to conduct layer-by-layer expansions between observed signals and hidden periodic states. Similarly, DEWP [59] is a also stack-by-stack expansion model to handle multivariate time series data, where each stack consists of a variable expansion block to capture dependencies among multiple variables and a time expansion block to learn temporal dependencies.

### v2

@@ Abstract
Time series, characterized by a sequence of data points organized in a discrete-time order, are ubiquitous in real-world scenarios. Unlike other data modalities, time series present unique challenges due to their intricate and dynamic nature, including the entanglement of nonlinear patterns and time-variant trends. Analyzing such data is of great significance in practical applications and has been extensively studied for centuries. Recent years have witnessed remarkable breakthroughs in the time series community, with techniques shifting from traditional statistical methods to contemporary deep learning models. In this paper, we delve into the design of deep time series models across various analysis tasks and review the existing literature from two perspectives: basic modules and model architectures. Further, we develop and release Time Series Library (TSLib) as a fair benchmark of deep time series models for diverse analysis tasks. TSLib implements 30 prominent models, covers 30 datasets from different domains, and supports five prevalent analysis tasks. Based on TSLib, we thoroughly evaluate 13 advanced deep time series models across diverse tasks. Empirical results indicate that models with specific structures are well-suited for distinct analytical tasks, providing insights for research and adoption of deep time series models. Code and datasets are available at https://github.com/thuml/Time-Series-Library.
@@
Time series refers to a sequence of data points indexed in a discrete-time order [1, 2], which are omnipresent in real-world applications, such as financial risk assessment, energy sustainability, and weather forecasting. Driven by the increasing availability of vast amounts of time series data across various domains, the community of time series analysis has witnessed tremendous advancements. Compared to image and text data, which have objectively prescribed syntax or intuitive patterns, the semantic information of time series data is primarily derived from the temporal variation [3]. This presents significant challenges in understanding the data, notably in identifying sequential dependencies, trends, seasonal patterns, and complicated dynamics. Consequently, analyzing time series data necessitates sophisticated methods to extract its meaningful temporal representations.
@@
Time series analysis has long stood as a pivotal research area, owing to its indispensable role in real-world applications [4, 5, 6]. It involves scrutinizing temporal variations in sequential data to uncover underlying patterns, thereby facilitating accurate predictions and informed decision-making. A fundamental aspect of this process lies in identifying intricate temporal dependencies and variable correlations inherent in the data. By capturing these entangled dependencies, time series models reveal underlying dynamics and support a broad spectrum of downstream tasks, including forecasting, classification, imputation, and anomaly detection.
@@
Traditional approaches, such as Autoregressive Integrated Moving Average (ARIMA) [1], Exponential Smoothing, and Spectral Analysis [7], have long served as stalwart tools for analyzing time series data. Grounded in statistical methodologies, these methods have been instrumental in uncovering patterns, trends, and seasonality within temporal variations. However, their rigid assumptions of linearity and stationarity limit their application to dynamic and evolving data flows, making them inadequate for capturing the intricate nonlinear relationships and long-term dependencies present in real-world time series data.
| Survey |
Analysis Task |
Model Architecture |
Benchmark | |||||||
Forecasting |
Classification |
Imputation |
Anomaly Detection |
MLP |
CNN |
RNN |
GNN |
Transformer |
Fawaz et al. (2019) [8] |
Braei et al. (2020) [9] |
Torres et al. (2021) |
García et al. (2021) |
Wen et al. (2022) [12] |
Jin et al. (2023) [13] |
Shao et al. (2023) [14] |
Qiu et al. (2024) [15] |
This Survey |
@@
Deep models have garnered wide attention and achieved remarkable performance across various domains, including natural language processing (NLP) [16], [17], computer vision (CV) [18], [19], and recommendation systems [20]. In recent years, deep learning models [21, 22, 23, 24, 3] have proven highly effective in capturing the intricate dependencies within time series data, making them a powerful tool for time series analysis over traditional statistical methods. More recently, Transformer models with attention mechanisms, originally developed for natural language processing tasks, have exhibited unprecedented capability in processing large-scale data [17] and have also been adapted for learning time series data. These architectures offer the advantage of selectively focusing on different parts of the input sequence, allowing for more nuanced discovery of both temporal and variable dependencies in time series.
@@
Although various time series models designed for different analysis tasks have emerged in recent years, there is a lack of a comprehensive overview of existing methods, covering both tasks and models. Previous reviews focus exclusively on either a specific model architecture or an analysis task. For example, [8, 9, 10, 11] each review deep learning methods for specific time series analysis tasks while failing to include advanced architecture such as Transformers. Several surveys [12, 13] provide up-to-date reviews for time series analysis focusing on specific deep learning architectures (e.g., Graph Neural Networks and Transformers). Recently, BasicTS [14] and TFB [15] introduce forecasting benchmarks that enable an unbiased evaluation of existing approaches but do not provide an overview of the architectural design of those deep models.
@@
In this survey, we provide a comprehensive review of deep time series models for researchers and practitioners, starting from the basic modules to modern architectures. To foster practical applications, a time series benchmark is offered for a fair evaluation and to identify the effective scope of existing models. This survey is organized as follows. Section 2 provides the background concepts of time series analysis. Section 3 introduces the basic modules that are widely utilized in prevalent deep time series models. Section 4 reviews existing deep time series models in terms of the architecture design. Section 5 introduces an easy-to-use open-source benchmark, Time Series Library (TSLib), along with extensive experimental comparisons and detailed analyses. Section 6 provides a brief discussion of future research directions while Section 7 summarizes this survey.
@@ 2.1 Time Series
Time series is a sequence of observations ordered by time, which can be denoted as , where represents the observed values at timestamp and is the number of variables. Note that spatiotemporal data (e.g. videos or weather records) can also be viewed as a special type of time series, where the “variate dimension” of the above notion corresponds to spatial pixels [25] in a video. However, previous research on spatiotemporal data primarily relies on highly structured spatial information [26], which limits their applications to general variates, like heterogeneous variates with distinct semantic meanings. In this paper, we focus on the more generic formalization, where we do not assume any spatial structure inside the variate dimension. Under this definition, the essence of time series analysis is to capture and utilize the temporal dependencies and inter-variable correlations within the observations.
Temporal Dependency Given the inherent sequential nature of time series data, one evident technological paradigm is to capture the temporal dependence of a set of historical data. Essentially, temporal dependencies refer to the intricate connections that manifest among distinct time points or sub-series. Traditional statistical methods, such as those based on autoregressive principles represented by ARIMA [1], have been widely studied to model these dependencies. Owing to their simplicity and interpretability, these statistical methods retain prominence in scenarios where the underlying temporal dynamics exhibit limited complexity.
Variate Correlation Beyond temporal dependencies, comprehending the correlations among distinct variables is paramount for analyzing multivariate time series, particularly in high-dimensional contexts. These correlations encapsulate the complex interactions and associations among multiple measurements as they evolve over time. Elucidating these relationships provides valuable insights into the underlying dynamics and latent processes. Conventional approaches, exemplified by VAR [27], extend the concept of autoregression to multiple variables, thereby enabling the capture of inter-variable relationships across time. However, VAR models represent each variable as a linear combination of lagged values of all other variables, which inherently limits their ability to capture misaligned or nonlinear relationships.
The aforementioned fundamental elements of time series present considerable challenges for constructing practical analysis models. While traditional statistical methods offer foundational theoretical insights, their inherent reliance on presumptions of the data, such as linearity and stationarity, often restricts their applicability in intricate real-world scenarios. Recently, deep learning models have emerged as powerful alternatives for time series modeling. Their capacity for representation learning and scalability to big data uniquely positions them for high-dimensional and nonlinear time series data, effectively addressing diverse analysis tasks.
@@ 2.2 Time Series Analysis Tasks
Based on the understanding of underlying patterns and trends within time series data, time series analysis encompasses various downstream applications, including forecasting [28, 29], imputation [30, 31, 32], classification [8, 33], and anomaly detection [34, 5], each finding distinct applications in diverse domains, as illustrated in Figure 1.
Forecasting is a fundamental task in time series analysis that requires models to uncover temporal dependencies and dynamic patterns. By capturing the relationships between past and future data, the forecasting model aims to predict future values or trends of the input series. Missing data due to sensor failures, data corruption, or absent measurements is ubiquitous in practical applications, leading to a growing demand for time series imputation to obtain higher-quality data. Unlike forecasting, which predicts future values based on historical observations, imputation focuses on reconstructing missing values using the available contextual information. Anomaly detection involves identifying unusual or abnormal patterns within a time series, which can indicate critical events, system faults, or outliers requiring further investigation. Lastly, classification assigns a label or category to a given time series based on its characteristics, a task widely utilized in fields such as medical diagnosis.
Given the pluralism of time series analysis, an array of deep learning models have emerged to address specific challenges in different tasks. Instead of a task-centric view into exhaustive details, this survey takes a model-centric view of deep time series models, focusing on their basic modules and architectural designs for addressing the foundational problems of capturing both temporal and inter-variable dependencies underlying large-scale time series datasets.
@@ 3 Basic Modules
In the past decade, time series modeling approaches have evolved significantly, shifting from traditional statistical models to sophisticated deep learning models. Despite these advancements, many classical tools and analytical algorithms remain widely used and continue to serve as foundational design principles in modern deep models. In this section, we focus on the major tools of classical time series analysis and demonstrate how they have been integrated as fundamental components in contemporary deep time series models.
@@ 3.1 Stationarization
Stationarity is a foundational concept in time series analysis, referring to the property where the statistical properties of a time series remain constant over time. A stationary time series exhibits a constant mean and variance, which simplifies statistical analysis and makes it easier to capture the underlying patterns and behaviors within a time series. Given that many statistics-based time series analysis methods fundamentally presuppose stationarity, the transformation of non-stationary data into a stationary form has become an essential preprocessing step. Conventionally, this stationarization is achieved through methods such as differencing or log-transformation. In recent deep learning approaches, data normalization [35] takes the role of stationarization in a simple but effective way, which standardizes the value distribution of observations while maintaining the intrinsic variations and further helps mitigate the distribution shift between the source and target domains.
Recent deep learning architectures have developed specialized layers to address the non-stationary problem. The deep adaptive input normalization (DAIN) layer [36] was proposed to adaptively standardize time series data based on its original distribution. RevIN [37] introduces a reversible instance normalization technique with learnable affine transforms to make the model bypass the non-stationary inputs. Further, Non-Stationary Transformer [38] (Stationary for short in the following) proposes a simpler but more effective series stationarization technique that improves the predictive capability of non-stationary series without extra parameters. Specifically, for a sequence with time stamps and variates , where , the outline of Stationary [38] can be summarized as follows:
where is a small value for numerical stability. are the variate-specific mean and variance. To recover the distribution and non-stationarity of the original series, a denormalization module is further used to augment the model output with mean and variance statistics of inputs.
The idea of stationarization and the aforementioned techniques have seen widespread adoption in subsequent deep time series models [23, 39, 40, 41]. For instance, recent works such as SAN [42] rethink the nature of non-stationary data by normalizing non-overlapping, equally-sized slices instead of the entire series, and incorporating a statistics prediction module to forecast future slice distributions.
@@ 3.2 Decomposition
Decomposition [43, 44], as a conventional approach in time series analysis, breaks time series into several components of more regular patterns, with a divide-and-conquer paradigm to explore complex series variations. In the previous work, a set of typical decomposition schemes are explored.
@@ 3.2.1 Seasonal-Trend Decomposition
Seasonal-trend decomposition [45, 46] is one of the most common practices to enhance the predictability of raw time series data. This technique separates a series into distinct additive components: trend, seasonal, cyclical, and irregular. The trend component represents the overall long-term pattern of the data over time, the cyclical component reflects repeated but non-periodic fluctuations within the data, the seasonal component indicates the repetitive patterns over a fixed period, and the irregular component is the residual or remainder of the time series after the other components have been removed.
Such decomposition can be achieved using mathematical tools like filters or exponential smoothing [47, 48]. Previous statistical approaches mainly adopt the trend-seasonality decomposition as data pre-processing [49]. In deep models, Autoformer [22] firstly introduces the idea of decomposition to deep learning architecture and proposes a series decomposition block as a basic module based on the average pooling to extract seasonal and trend-cyclical parts of deep features and input series:
Implemented based on a temporally average pooling layer, the proposed decomposition block effectively captures the trends , and the remainder is taken as the seasonal part . This design has been widely adopted as a native building block in the subsequent works [6, 50, 51, 52, 53] for disentangling underlying patterns within deep features.
@@ 3.2.2 Basis Expansion
Basis expansion is a mathematical method used to represent a function or a set of data points in terms of a new set of pre-defined functions. These new functions form the bases for a function space, allowing any function in that space to be expressed as a linear combination of these basis functions. In the context of time series analysis, basis expansion is used to reveal nonlinear temporal relationships by decomposing the time series into a combination of basic variations, thereby enhancing interpretability. As a representative model, N-BEATS [54] presents hierarchical decomposition to time series by utilizing a fully connected layer to produce expansion coefficients for both backward and forward forecasts. For the -th block in the proposed hierarchical architecture, the operation is
@@
Subsequently, N-HiTs [55] redefine the N-BEATS by incorporating subsampling layers before the fully connected blocks. This modification enhances input decomposition through multi-frequency data sampling and improves future prediction via multi-scale interpolation. DEPTS [56] proposes a novel decoupled formulation for periodic time series by introducing a periodic state as a hidden variable. It then develops a deep expansion module atop residual learning to perform layer-by-layer expansions between observed signals and these hidden periodic states. Similarly, DEWP [57] is a stack-by-stack expansion model designed to handle multivariate time series data. Each stack within the model comprises a variable expansion block for capturing dependencies among multiple variables and a time expansion block for learning temporal dependencies.